OpenStudy (anonymous):

1. P=2L+2w p=700 meters w= .75 x L 700/4 sides 175 =4 equal sides i substituted 150 in for w. then subed 200 in for L L= 200 W= 150 2.The width of my computer room is _____% of its length. If the perimeter is 52 feet, then what are the length and width?

OpenStudy (anonymous):

TIANA

OpenStudy (amistre64):

can you format that to something readable by the rest of us :)

OpenStudy (anonymous):

it is a discussion topic i have to answer that is what is given

OpenStudy (anonymous):

1. Solve the following problem. The width of a rectangular lot is 75% of its length. If the perimeter is 700 meters, then what are the length and width?

OpenStudy (amistre64):

P=2l+2w is the perimeter of a rectangular room. P=700 width is 3/4 of the lenght...

OpenStudy (amistre64):

700 = 2l + 6l/4

OpenStudy (anonymous):

I'm not sure what the question is in the first one...because i think you solved it... for the second question let L = length let W = width let x = percentage (ie 5%, then x would be 0.05) W = L*x 2W + 2L = 52 substitute 2(L*x) + 2L = 52 L*x + L = 26 L ( x + 1) = 26 L = 26 / (x+1) W = (26*x) / (x+1)

OpenStudy (amistre64):

700 = 14l/4 280 = 14l 280/14 = l l = 140/7 = 120

OpenStudy (anonymous):

tina

OpenStudy (amistre64):

l = 700; w = 350/4 if we go the top route and thats after I fixed my glaring fauxpas :)

OpenStudy (amistre64):

DOH!.....

OpenStudy (amistre64):

i think i just had a lobotomy :)

OpenStudy (amistre64):

l=200; w=150

OpenStudy (amistre64):

whats with number 2?

OpenStudy (anonymous):

# 2 is me having to measure a room and make up an equation for someone else to solve or do you mean the 2L 2W

OpenStudy (amistre64):

hmmm... then say the length is 100% of the width and make it a square ;)

OpenStudy (amistre64):

52/4 = 13....that was EASY!! lol

OpenStudy (anonymous):

amistre plz helo me write that essay its do tommorwo

OpenStudy (anonymous):

cant be a square has to be rectangluar

OpenStudy (amistre64):

a square IS a rectangle

OpenStudy (anonymous):

amistre plz

OpenStudy (amistre64):

by definition a square is a rectangle; just a special rectangle

OpenStudy (anonymous):

ok so what is the answer you confused me

OpenStudy (amistre64):

mary, I dont write esssays; at least not well. And I aint got the time to do no reasearchand put the stuff together if I did.

OpenStudy (amistre64):

You wanted to make up a question to fit the ...thing.. in Q2 right?

OpenStudy (amistre64):

if p=52 and the width is 100% of the l then l and w are 13

OpenStudy (anonymous):

Q1 and Q2 are for my discussion thread it belongs to another student that I have to answer, Pt.2 is i have to do the same for him to answer

OpenStudy (amistre64):

then put in 100% as the answer :) "The width of my computer room is __100___% of its length. If the perimeter is 52 feet, then what are the length and width?"

OpenStudy (amistre64):

l=13 ; w=13

OpenStudy (anonymous):

ok gottcha

OpenStudy (anonymous):

1. Solve the following problem. The width of a rectangular lot is 75% of its length. If the perimeter is 700 meters, then what are the length and width? 2. Measure the length and width of any rectangular (not square) room in your residence, and calculate the perimeter. Determine what percent of the length the width of the room is. Write a problem following the format below and post it for your classmates to solve. The width of my (living room, family room, etc) is ____% of its length. If the perimeter is ____ meters, then what are the length and width? Due by 11:59 MST on Friday. 3 Pick a problem posted by your classmate and solve it either by substitution or elimination. Make sure to pick the problem not solved by somebody else and show all steps in the solution. Explain why you chose the method you did. Due by 11:59 MST on Sunday.

OpenStudy (anonymous):

i know thats alot but this is the assignment so u helped me answer the students which is number 3

OpenStudy (amistre64):

percent is figured out by taking the "part" and dividing by the "total"; times 100. 35/100 = .35 .35 * 100 = 35% :)

OpenStudy (amistre64):

if the length is 144 and the width is 98: the percentages are: l/w = .xxx times 100 = xx.x%

OpenStudy (anonymous):

what is #1

OpenStudy (amistre64):

L=200; W=150 is #1

OpenStudy (anonymous):

ok i have to show my work

OpenStudy (amistre64):

W = (.75)L 700 = 2(.75L) + 2L 700 = 1.5L +2L 700 = 3.5L 700/3.5 = L L = 200. W = .75(200) = 150

OpenStudy (amistre64):

so number 2 says NOT a square...what a crock !

OpenStudy (amistre64):

pick a number....

OpenStudy (amistre64):

between 1 and 15

OpenStudy (anonymous):

7

OpenStudy (anonymous):

does 7 work

OpenStudy (anonymous):

r u there

OpenStudy (amistre64):

im here.. :) got side tracked

OpenStudy (amistre64):

....butterfly......>>>>>

OpenStudy (anonymous):

i love butterflies all my tats are

OpenStudy (amistre64):

7 is fine; say 7 is the width and times that by 4; 28 28 is my length; and my width is 25% of the length

OpenStudy (amistre64):

14+56 = 70 is the perimeter of my room....

OpenStudy (amistre64):

i aint got any tats. im a good Mormon, most of the time :)

OpenStudy (anonymous):

ive got 3

OpenStudy (anonymous):

The width of my room, is 25% of its length. If the perimeter is ____ meters, then what are the length and width? the perimeter is how many meters 70?

OpenStudy (amistre64):

70..yes/ 7+7+28+28 = 14+56 = 70

OpenStudy (anonymous):

ok i get it

OpenStudy (anonymous):

hey r u there amistre64? yesterday you helped me with this The width of my room, is 25% of its length. If the perimeter is 70 meters, then what are the length and width? was that 70meters or 70 ft this is my teacher response Please check your measurements. First, the ratio of the width to the length is too small. Your room must very narrow. Second, the perimeter of 70 meters appears to be to big for a room in a residence. Is it in meters or feet?